A candy factory has a large vat into which workers pour chocolate and cream. Each ingredient flows into the vat from its own special hose, and each ingredient comes out of its hose at a constant rate. Workers at the fatory know that it takes 20 minutes to fill the vat with chocolate from the chocolate hose, and it takes 15 minutes to fill the vat with cream from the cream hose. If workers pour both chocolate and cream into the vat at the same time (each coming full tilt out of its own hose). how long will it take to fill the vat? Before you find an exact answer to this problem find an approximate answer or find a range, such as "between and ...minutes," Explain your answer.
I also need to know the name of the strategy used to solve the problem.

Answer by stanbon(57387)

Workers at the factory know that it takes 20 minutes to fill the vat with chocolate from the chocolate hose, and it takes 15 minutes to fill the vat with cream from the cream hose.
------
Chocolate hose rate: 1/20 job/min
Cream hose rate: 1/15 job/min
----
Together rate: 1/x job/min
---------------------------------
Equation:
rate + rate = together rate
1/20 + 1/15 = 1/x
-----------------
Multiply thru by 60x to get:
3x + 4x = 60
7x = 60
x = 60/7 minutes = 8 4/7 minutes (time to fill the vat together)
============================
Cheers,
Stan H.